OpenStudy (anonymous):

A population of 100 fish in an aquarium consists of 49 homozygous dominant green individuals, 42 heterozygous green individuals, and 9 white individuals. What are the frequencies of the green and white alleles in the gene pool? PLEASE I NEED HELP UNDERSTANDING! medals given.

OpenStudy (anonymous):

The frequency of GG is 0.49, so the frequency of G is sqrt(0.49) = 0.7. Therefore, the frequency of g is 1-0.7 = 0.3. The frequency of GG is (0.7)↑2 = 0.49, of Gg is 2(0.7)(0.3) = 0.42, and of gg is (0.3)↑2 = 0.9. If the population is in equilibrium, the frequencies will remain the same.

OpenStudy (anonymous):

GG=green? gg=white? so white is 30% and green is 70%?

OpenStudy (anonymous):

yes i believe

OpenStudy (anonymous):

what are you writing mrdoldum?

OpenStudy (anonymous):

@andrewjordan833 So long as the many requirements of the Hardy-Weinberg's many requirements the frequencies will be the same. No selection, infinite population, completely random mating, no mutation, no migration (gene flow) etc etc. @jyar. This is a dom/ recessive trait, that means we have two alleles of one gene. One way: Let's set the dominant allele as G and the recessive as g. We know that the total frequencies of these two alleles must be 1. Why? Frequency is the same as saying of the total number of G and g alleles, what percentage is G? or g? 49 GG 42 Gg 9 gg Now when we find the totals we have \[49*2+42=140\] So, there are 140 G alleles \[9*2+42=60\]So, 60 g alleles. There is a total of 200 alleles (includes both G and g). Now we see the frequencies are:\[140/200=0.7\] Way two, using ol' hardy: p=frequency of G and q=frequency of g \[p^{2}+2pq+q^{2}=1\] Here, \[p^{2}\] is the 0.49 so just \[p\] is the frequency of the G allele and \[\sqrt{0.49}=frequency G=0.7\]

OpenStudy (anonymous):

i would go with @mrdoldum explanation

OpenStudy (anonymous):

thank you @mrdoldum thats amazing!